OpenStudy (anonymous):

A train leaves watsonville at noon traveling north at 44 mph. one hour later, another train, going 52 mph, travels north on a parallel tack. how many hours will the second train travel before it overtakes the first train. Show ur step please

OpenStudy (dumbcow):

5.5 hours

OpenStudy (anonymous):

can i get you to show ur step please...

OpenStudy (dumbcow):

d=r*t when train 2 overtakes train 1 they both will have gone same distance with 1 hour time difference 44(t+1) = 52t 44t +44 = 52t -8t =-44 t = 44/8 t=11/2